OpenStudy (scarlettthecamel):

q is inversely proportional to the square of r. q=4 when r=5 (A) Find the formula for q in terms of r (B) Work out the value of q when r=15

OpenStudy (mathstudent55):

y is directly proportional to x: \(y = kx\) y is inversely proportional to x: \(y = \dfrac{k}{x}\)

OpenStudy (mathstudent55):

In your case, start with the second formula above. If you had q is inversely proportional to r, you'd have \(q = \dfrac{k}{r} \), but you have q is inversely proportional to the "square of r," so we much account for the "square of r" part: \(q = \dfrac{k}{r^2} \)

OpenStudy (scarlettthecamel):

okkkkkkkkkkk

OpenStudy (mathstudent55):

Now you need to find k. To do that, use the last equation above, and enter the known point, q = 4 when r = 5, and find k. That answers part (A). For part (B), use the formula with the value of k you find to find the value of q when r = 15.

OpenStudy (scarlettthecamel):

Could you go into more detail about finding k?

OpenStudy (mathstudent55):

Sure. Start with \(q = \dfrac{k}{r^2}\) Now substitute into that equation the given values of r and q. \(4 = \dfrac{k}{5^2}\) Now can you solve for k?

OpenStudy (scarlettthecamel):

do i do opposite to divide so times?

OpenStudy (mathstudent55):

Square the 5 first. The multiply both sides by the square of 5.

OpenStudy (mathstudent55):

What is 5^2 = ?

OpenStudy (scarlettthecamel):

ok thanks!!!!!!!

OpenStudy (scarlettthecamel):

25

OpenStudy (scarlettthecamel):

so 4 times 25?

OpenStudy (mathstudent55):

Good. Now multiply both sides by 25 to get rid of the 25 in the denominator. Yes.

OpenStudy (mathstudent55):

k = 25 * 4

OpenStudy (scarlettthecamel):

thank you so much!! you have actually explained it so that i get it, better than my maths teacher lol

OpenStudy (mathstudent55):

You're welcome. I hope you got k = 100, right?

OpenStudy (scarlettthecamel):

ok is that what i write for answer a?

OpenStudy (scarlettthecamel):

k=100

OpenStudy (mathstudent55):

Now that you know what k is, you rewrite the formula \(q = \dfrac{k}{r^2} \) with 100 instead of k. That is the answer to part (A).

OpenStudy (scarlettthecamel):

ohhhh thanks!

OpenStudy (mathstudent55):

(A) \(q = \dfrac{100}{r^2} \)

OpenStudy (scarlettthecamel):

thankyou!!!!!!!!!!!!

OpenStudy (mathstudent55):

Now you know what the equation for the inverse proportion is. To solve part (B), just use that formula, and substitute r with 15, and solve for q.